A local hospital is holding a two day marathon walk to raise funds for a new research facility. The total distance of the marathon is 26.2 miles. On the first day, Martha starts walking at 10:00 am. She walks 4 miles per hour. Carol starts two hours later than Martha but decides to run to catch up to Martha. Carol runs at a speed of 6 miles per hour. Carol wants to reduce the time she takes to catch up to Martha by 1 hour. How can she do this by changing her starting time? How can she do this by changing her speed?
Answer by mananth(16946)

Martha --speed = 4mph---10.00 am
Carol----speed ---6mph---12.00 noon
x/4-2=x/6
x/4-x/6=2
(6x-4x)/24=2
2x=48
x=24 distance where she catches up
Carol distance =24
carol speed = 6
carol time = 4 hours
reduce by 1 hour
carol time will be 3 hours
distance = 24miles
speed = 24/3 = 8 mph